# Holodeck 2.0: Vision‑Language‑Guided 3D World Generation with Editing

### v1

@@ Introduction
Scene Generation
Spatial Constraint Generation. The Scene Generation Module first employs GPT‑o3 to produce a series of spatial constraints between objects based on the text input, the scene reference image, and object properties. We define a total of ten spatial relations, which are divided into the following four categories: (1) Relative: in front of, behind, left of, right of, side of; (2) Distance: near, far; (3) Vertical: on, above and (4) Rotation: face to. As shown in Figure 4, large reasoning models can generate realistic and spatially plausible constraints for objects.
Layout Generation. For layout generation, we employ a DFS solver to compute object placements based on object properties and spatial constraints between objects. We define mathematical formulations to enforce each spatial constraint (for example, “face to” means that the angle between the source object’s orientation and the line connecting the centers of the two objects is less than 10 degrees). Objects should satisfy these spatial constraints while avoiding physical collisions. To fully leverage GPT-o3’s spatial reasoning capabilities, we prioritize the initial object positions inferred by the model. It is worth noting that, when the DFS solver encounters spatial constraints that are difficult to satisfy simultaneously, the Scene Generation Module re‐prompts GPT‑o3 using the DFS solver’s feedback to update the objects’ spatial constraints. The DFS solver and the reasoning model then form an iterative loop until a valid layout is found. This approach significantly enhances Holodeck 2.0’s layout capabilities. We present several examples of position plots for the bounding boxes produced by the solver (Figure 4).
Scene Editing
Scene Layout Level. If users wish to personalize the generated scene layout after it has been created, they can provide a natural-language edit instruction. Then, it is incorporated into the original workflow to regenerate the spatial constraints between objects, and the DFS solver iteratively computes the updated layout.
3D Asset level. Utilizing 3D generative models, we support asset‑level editing. Because our scenes are built by assembling individual 3D assets, we can easily delete objects in response to user language feedback. Users can also replace objects, for example, “Change the torii gate to a pastel gradient color scheme and hang decorative bells on it.” The Scene Editing Module then uses that language feedback to regenerate a style‑appropriate 3D asset for the specified object and substitute it into the scene. When adding new objects, the module regenerates suitable 3D assets and recalculates spatial constraints to maintain coherent scene layouts.
Applications in Gaming
We explore the potential of directly integrating Holodeck 2.0-generated scenes into commercial game engines for game development. To demonstrate this, we use Holodeck 2.0 to construct a detailed museum exhibition hall based on rich textual descriptions specifying each artifact’s type, material, style, and placement. Holodeck 2.0 produces a 3D museum scene that closely matches the text, featuring the Temple of Dendur at its center, surrounded by over a dozen objects such as the Venus de Milo and a Shang‑dynasty bronze ding cauldron. This generated 3D scene can then be directly imported into professional game engines such as Unreal Engine, enabling quick and intuitive development of interactions between the game character and artifacts. Such a virtual museum game can serve both educational and entertainment purposes. The results in Figure 6 demonstrate Holodeck 2.0’s high-quality, highly faithful scene generation from detailed text, highlighting its potential for direct application in game content creation.
Evaluation
To systematically evaluate the quality of 3D scenes generated by Holodeck 2.0, we employ a combination of human evaluations (involving more than 300 responses, the majority of which are from undergraduate and graduate students) and CLIP‑based automated scoring. Holodeck (Yang et al. 2024) is a representative method for 3D scene generation that retrieves objects from 3D asset libraries and is one of the most commonly used baselines in recent work. Although it primarily focuses on indoor scene generation, its open‑vocabulary text input makes it applicable to open‑domain scenes to some extent. Therefore, we choose to directly compare with the Holodeck method. Specifically, we generated 60 test cases spanning both indoor and open‑domain scenes. Our comparisons show that Holodeck 2.0 outperforms the baseline Holodeck in both indoor and open‑domain settings, with the advantage being more pronounced in open‑domain scenarios. In addition, we perform ablation studies on the Scene Generation Module. We further compare the Scene Editing Module with the baseline method, BlenderAlchemy (Huang, Yang, and Guibas 2024). The detailed results are shown below.
Comparative Analysis
Setup. We use LLMs to generate 30 input texts as test cases for indoor scenes and another 30 for open-domain scenes (see the supplement for examples of detailed descriptions, each averaging about 50 words). These indoor scenes span different categories, including bedrooms, gyms, playrooms, museums, and laboratories. These open-domain scenes are varied, including architectural complexes, natural landscapes, and futuristic environments. They also exhibit different styles, such as realistic, cartoon and steampunk. We then feed each text into both Holodeck and Holodeck 2.0 to produce 60 paired scenes (30 paired indoor scenes and 30 paired open-domain scenes) for human evaluation and CLIP‑based scoring. In the human evaluation experiment, we render each 3D scene from the least occluded viewpoint using identical parameters, then present the rendered images to annotators in shuffled order within each pair. Annotators score each pair on a scale from 1 to 10 along two dimensions: (1) Asset level: the overall quality of the 3D assets displayed in the scene; (2) Scene level: the overall quality of the scene with respect to the input description. For CLIP-based scoring, we measure the embedding similarity (CLIP ViT-B/32) between each generated scene’s rendered image and the input text to evaluate the quality of the generation.
Methods Asset-level Scene-level CLIP Score
Holodeck 4.28 3.67 0.253
Holodeck 2.0 8.00 7.71 0.299
Table 1: Evaluation results on Indoor Scenes. We have human-evaluated asset-level and scene-level scores, as well as the automatically computed CLIP Score.
Methods Asset-level Scene-level CLIP Score
Holodeck 3.77 3.06 0.256
Holodeck 2.0 7.58 7.32 0.307
Table 2: Evaluation results on Open-domain Scenes.
Results. Each of the indoor and open-domain scene evaluation surveys is completed by 105 annotators. Table 1 and Table 2 present the quantitative scores (human evaluation and CLIP-based scoring) for indoor scenes and open-domain scenes, comparing Holodeck 2.0 and Holodeck. Figure 7 illustrates the full 1–10 rating distributions from the human evaluation, showing how scores are spread across both asset and scene levels. Figure 8 shows the CLIP score distributions. These results show Holodeck 2.0 outperforms Holodeck in both indoor and open‑domain settings, with the advantage being more significant in the latter, demonstrating its strong capability in scene generation tasks.
Figure 7: Human evaluation on 1–10 rating distributions at asset and scene levels for indoor (left) and open-domain (right) scenes, comparing baseline Holodeck (blue hues) versus Holodeck 2.0 (red hues).
Figure 8: Comparison of CLIP score distributions for Holodeck (baseline) versus Holodeck 2.0 (ours). Kernel density estimates are shown for 30 indoor scenes (left) and 30 open-domain scenes (right), with dashed lines marking the mean score of each method.
Figure 9: Comparative human evaluation of Holodeck 2.0 across two experiments. Pie charts display annotator preference percentages and counts for Ablation Study (left) and Scene Editing Comparison (right).
Ablation Study on the Scene Generation Module
Setup. We also conduct an ablation study on Holodeck 2.0’s Scene Generation Module. We evaluate two configurations: (1) Using only the initial layouts from the VLM; (2) Using the full Scene Generation Module. For each input text from the Comparative Analysis, we generate 2 scenes (with the same 3D assets) under both settings and then assess their quality using human evaluation. We asked human evaluators to choose among 60 pairs of scene layouts, with three options available: “Prefer A”, “Prefer B”, and “Tie”.
Results. 94 annotators complete the survey, with each annotator answering all 60 questions. Human evaluation shows that the full Scene Generation Module is indispensable for effective 3D scene generation. As the left panel of Figure 9 shows, 67.93% of the choices favor the scenes after our layout, 12.89% prefer the initial placements generated directly by the VLM, and the remaining 19.18% are ties.
Evaluation on Scene Editing
Previous 3D scene editing methods, such as BlenderGPT and BlenderAlchemy (Huang, Yang, and Guibas 2024), have mainly focused on material design and lighting systems, but struggle with style replacement and object placement. By fully utilizing 3D asset generation models, we accomplish what methods like BlenderAlchemy struggle to achieve, such as adding objects in specific artistic styles or replacing existing objects with customized stylistic variations.
For object placement editing, since BlenderAlchemy can achieve this by editing Blender scripts with VLMs, we conduct a human evaluation to compare Holodeck 2.0 against BlenderAlchemy in placement editing tasks. In our control experiment: (1) We perform 10 representative placement editing tests for both BlenderAlchemy and Holodeck 2.0. (2) Annotators are then required to choose their preferred version between the two edited scenes.
@@
197 annotators complete the survey, with each annotator answering all 10 questions. As the right panel of Figure 9 shows, 71.12% of annotators prefer our edited scenes, while only 28.88% favor BlenderAlchemy’s outputs. This demonstrates that our integrated approach combining DFS and VLM-generated spatial constraints for layout editing outperforms the method relying solely on VLM-based editing and evaluation.
Conclusions and Discussions
@@
We propose Holodeck 2.0, a vision-language-guided system, equipped with 3D asset generation models, that can generate and edit diverse, customizable, and stylistically varied 3D worlds. We validate the effectiveness of Holodeck 2.0 through human evaluations and CLIP-based scoring. We also show its potential in game modeling. In future work, we intend to improve Holodeck 2.0’s capacity for generating more complex and finely detailed 3D worlds, thereby facilitating the direct creation of large-scale game environments.
Acknowledgements
We thank Ben Swanson for his help in improving our understanding of 3D generative models and game engines, and for providing helpful suggestions on the manuscript.
Appendix
Details of Holodeck 2.0
Details of the Scene Analysis Module
@@ 12 I have the following list of image filenames:
Details of the Object Generation Module
In the Object Generation Module, we run the deployed Hunyuan3D 2.1 model locally. Given the individual image of each object, Hunyuan3D 2.1 generates a corresponding 3D asset in GLB-format. Since generating each 3D asset takes approximately 3–5 minutes on an A6000 GPU, Holodeck 2.0 parallelizes the asset generation process based on available resources. After generating the 3D asset for each object, we write the asset’s file path back into the JSON file containing the object properties. We also rescale each 3D asset based on its actual x, y, and z dimensions, using the original z size as a reference. The updated x and y sizes after scaling are then written back into the JSON file.
Details of the Scene Generation Module
@@
Spatial Constraints Definition
@@
[…]
@@
Details of the Scene Editing Module
@@
About the Input Type
Holodeck 2.0 is capable of generating scenes from finely detailed textual descriptions, producing results that closely match the input. However, we also recognize that in some cases, users may not have highly specific requirements and may provide only a brief, high-level description. In such scenarios, we can leverage an additional LLM to expand the user’s simple input into a more detailed scene description, which can then be used to generate the corresponding scene.
Holodeck 2.0 also supports generating a scene directly from a single image, effectively using the scene reference image as the starting point for scene generation.
Rendering Options
@@
Qualitative Examples
Details of Evaluation
We conduct a total of four online questionnaires; the first two focus on static scene generation in indoor and outdoor environments, respectively. Each of these two questionnaires comprises 30 distinct scenarios. For each scenario, participants answer four separate rating questions on a 10-point Likert scale (1 = worst, 10 = best), measuring:
  • Object‐level quality of Holodeck v.s. Holodeck 2.0
  • Scene‐level quality of Holodeck v.s. Holodeck 2.0
@@
The two rendered 3D scenes generated from this text are shown below.
(a) Rendered Image of 3D Scene A
(b) Rendered Image of 3D Scene B
@@
The two rendered 3D scenes generated from this text are shown below.
(a) Rendered Image of 3D Scene A
(b) Rendered Image of 3D Scene B

### v2

@@ 3.4 Scene Editing
4 Experiments
To systematically evaluate the quality of 3D scenes generated by Holodeck 2.0, we employ a combination of human evaluations, CLIP‑based scoring and VLM-based scoring in Section 4.1. Specifically, we generate 70 test cases spanning both indoor and open‑domain scenes. Our comparisons show that Holodeck 2.0 outperforms the baselines in both indoor and open‑domain settings, with the advantage being more pronounced in open‑domain scenarios. In Section 4.2, we perform ablation studies on the Scene Generation Module. In Section 4.3, we further compare the Scene Editing Module with the baseline methods, BlenderAlchemy [huang2024blenderalchemy] and BlenderGym [gu2025blendergym]. In Section 4.4, we additionally demonstrate Holodeck 2.0’s practical applicability through direct integration with commercial game engines. The detailed results are shown below.
Table 1: Benchmarking performance on indoor and open-domain scenes. We report perceptual, physical, and spatial-semantic metrics. GraphDreamer’s metrics are limited due to its implicit representation (see supplement).
Indoor Open-domain
Method Perceptual Physical Spatial-Semantic Perceptual Physical Spatial-Semantic
[…]
Holodeck 0.256 4.70 3.99 51.43 1.89 67.83 69.06 38.04 0.257 5.55 4.80 37.14 3.51 72.29 68.86 27.67
PartCrafter 0.259 5.02 4.55 71.42 0.29 80.03 78.91 58.76 0.231 3.09 2.43 100.00 0.00 57.06 61.11 64.00
Scene Language 0.210 2.81 2.76 47.37 1.53 74.09 65.73 37.06 0.216 2.19 2.10 36.00 2.56 65.33 71.15 27.71
GraphDreamer 0.170 15.50 12.25 0.226 26.17 26.67
Ours 0.297 8.31 7.86 97.14 0.11 81.03 82.86 82.67 0.307 8.07 7.84 97.14 0.86 84.43 83.63 86.00
Table 2: Ablation study results. HP denotes human preference when comparing two scenes with different layouts: w/o Layout Solver and our full method. (Tied cases are split equally between both methods.)
Indoor Open-domain
Method Perceptual Physical Spatial-Semantic Perceptual Physical Spatial-Semantic
[…]
w/o Layout Solver 0.289 24.63 54.29 1.23 77.83 76.11 45.30 0.306 23.70 42.86 1.89 79.09 77.83 36.53
Ours 0.297 75.37 97.14 0.11 81.03 82.86 82.67 0.307 76.30 97.14 0.86 84.43 83.63 86.00
4.1 Comparative Analysis
Metrics. We evaluate generated 3D scenes along three complementary dimensions: perceptual, physical, and spatial-semantic. A detailed description of all metrics is provided below:
(1) Perceptual Quality
  • CLIP score: Text–image similarity between scene rendering and input text, computed using CLIP (ViT-B/32).
  • Human rating at the Asset level (HA): Human evaluation of individual 3D assets’ realism and fidelity.
  • Human rating at the Scene level (HS): Human evaluation of overall scene plausibility and consistency with the input text.
(2) Physical Plausibility
  • Collision-Free ratio (CF): The percentage of scenes without collisions between objects.
  • Collision Number (CN): Average number of collision pairs per scene.
(3) Spatial-Semantic Consistency
  • Positional Coherency (Pos.): Measures how well object positions match the spatial relations indicated by the text.
  • Rotational Coherency (Rot.): Measures how well object orientations match the directional relations indicated by the text.
  • Physically-grounded Semantic Alignment (PSA): A holistic metric that evaluates semantic spatial correctness together with physical feasibility.
The complete formulation of those metrics can be found in the supplement.
Baselines. We compare Holodeck 2.0 against four representative 3D scene generation approaches: Holodeck [yang2024Holodeck], PartCrafter [lin2025partcrafter], Scene Language [zhang2025scenelanguage], and GraphDreamer [gao2024graphdreamer]. Holodeck [yang2024Holodeck] is a representative method for 3D scene generation that retrieves objects from 3D asset libraries. PartCrafter [lin2025partcrafter] jointly generates multiple 3D parts and objects from a single image using compositional latent diffusion transformers. Scene Language [zhang2025scenelanguage] represents scenes with programs, words, and embeddings inferred from language models for 3D/4D generation. GraphDreamer [gao2024graphdreamer] takes a scene graph as input and generates a compositional 3D scene where each object is fully disentangled.
Setup. We construct a benchmark of 70 text inputs, comprising 35 indoor scenes and 35 open-domain scenes. (See the supplement for examples of detailed descriptions, ranging from 50–300 words). Indoor text inputs describe environments such as bedrooms, gyms, classrooms, and laboratories, while open-domain text inputs include plazas, city streets, forests, and theme-park settings.
To accommodate different input requirements across methods, we apply necessary preprocessing while maintaining fairness. Since PartCrafter [lin2025partcrafter] requires image inputs rather than text, we use GPT-Image-1 to convert text descriptions into reference images. For GraphDreamer [gao2024graphdreamer], which requires object-level descriptions as input, we use preprocessing scripts to extract and format them from the full scene descriptions.
For each text input, we generate corresponding scenes using all four baseline methods and Holodeck 2.0. Human evaluation is conducted with responses from undergraduate and graduate students, who rate each rendered scene on a 1–10 scale at both the asset and scene levels. Automated metrics (CLIP, CF, CN, Pos., Rot., PSA) are computed directly from the rendered outputs using a unified evaluation script across all methods.
Results. The indoor and open-domain surveys are completed by 183 and 268 annotators, respectively. Table 1 reports quantitative results across all metrics, including human assessments, the CLIP-based perceptual score, physical plausibility measures, and spatial-semantic alignment scores.
Overall, Holodeck 2.0 achieves the best performance across nearly all metrics in both indoor and open-domain settings. For perceptual quality, Holodeck 2.0 attains the highest CLIP score (0.297 for indoor and 0.307 for open-domain) and substantially higher human evaluation scores at both asset and scene levels. Physical plausibility also improves significantly, with Holodeck 2.0 maintaining a near-perfect CF (97.14%) while keeping the CN low (0.11 for indoor and 0.86 for open-domain). In terms of spatial-semantic coherence, Holodeck 2.0 yields the best positional and rotational consistency and the highest PSA scores (82.67 for indoor and 86.00 for open-domain), indicating precise alignment between generated layouts and textual spatial relations.
Among the baselines, PartCrafter performs the best after Holodeck 2.0, particularly in physical stability, while Holodeck and Scene Language show moderate perceptual alignment but suffer from frequent object collisions. GraphDreamer shows lower performance than others. Overall, these results demonstrate that Holodeck 2.0 achieves superior perceptual realism, physical validity, and semantic coherence across both indoor and open-domain scenes.
4.2 Ablate the Scene Generation Module
Setup. We also conduct an ablation study on Holodeck 2.0’s Scene Generation Module. We evaluate two configurations: (1) Using only the initial layouts from the VLM (without the Layout Solver); (2) Using the full Scene Generation Module. For each input text from the Comparative Analysis, we generate 2 scenes (with the same 3D assets) under both settings and then assess their quality using human evaluation. We ask human annotators to choose among 70 pairs of scene layouts, with three options available: “Prefer A”, “Prefer B”, and “Tie”. Other automated metrics, identical to those in Table 1, are also computed.
Results. A total of 145 annotators complete the survey. Human evaluation shows that the full Scene Generation Module is indispensable for effective 3D scene generation. As shown in Table 2, 75.37% of responses favor our full method for indoor scenes, while 76.30% favor it for open-domain scenes. Regarding physical plausibility, without the Layout Solver, the CF drops to roughly half of that of the full method. Spatial-semantic metrics also decrease significantly: the PSA drops from 82.67 to 45.30 for indoor and from 86.00 to 36.53 for open-domain settings.
@@ 4.2 Ablate the Scene Generation Module
Figure 7: Human evaluation results on scene editing. The rating distributions (1–10 scale) of Holodeck 2.0, BlenderAlchemy, and BlenderGym are shown for object placement tasks.
4.3 Evaluation on Scene Editing
Baselines. We compare Holodeck 2.0 against two representative 3D scene editing methods: BlenderAlchemy [huang2024blenderalchemy] and BlenderGym [gu2025blendergym]. BlenderAlchemy [huang2024blenderalchemy] leverages VLMs to generate Blender editing sequences by iteratively refining Python programs from user intentions. BlenderGym [gu2025blendergym] is an extension of BlenderAlchemy, with better performance in object placement tasks.
Setup. We conduct a human evaluation to compare Holodeck 2.0 against both BlenderAlchemy and BlenderGym in placement editing tasks. In our evaluation study: (1) We select 10 representative placement editing tests and apply all three methods to each, producing 30 edited scenes in total. (2) Annotators are then required to rate each edited scene on a scale of 1 to 10 based on overall editing quality and consistency with the instructions.
@@ 4.3 Evaluation on Scene Editing
Results. 450 annotators complete the survey. As shown in Figure 7, our edited scenes achieve the highest average rating of 9.06, compared to 4.29 for BlenderAlchemy and 4.87 for BlenderGym. This demonstrates that our integrated approach combining DFS and VLM-generated spatial constraints for layout editing outperforms the methods relying solely on VLM-based editing.
In addition to object placement tasks, our method also surpasses these two approaches in style replacement, handling tasks they struggle with, including adding objects in particular artistic styles and replacing existing objects with customized stylistic variations (as shown in Figure 5).
4.4 Applications in Gaming
Furthermore, we explore the potential of directly integrating Holodeck 2.0-generated scenes into commercial game engines for game development. To show this, we utilize Holodeck 2.0 to construct a detailed museum exhibition hall based on rich textual descriptions specifying each artifact’s type, material, style, and placement. Holodeck 2.0 produces a 3D museum scene that closely matches the text, featuring the Temple of Dendur at its center, surrounded by over a dozen objects such as the Venus de Milo and a Shang‑dynasty bronze ding cauldron. This generated 3D scene can then be directly imported into professional game engines such as Unreal Engine, enabling quick and intuitive development of interactions between the game character and artifacts. Additionally, physics is enabled to ensure realistic object interactions. The results in Figure 6 demonstrate Holodeck 2.0’s high-quality, highly faithful scene generation from detailed text, highlighting its potential for direct application in game content creation.
5 Conclusions and Discussions
@@
We propose Holodeck 2.0, a unified vision-language-guided framework that achieves high-fidelity, stylistically consistent, and editable 3D world generation. We validate the effectiveness of Holodeck 2.0 through human and model evaluations. We also show its potential in game modeling. In future work, we intend to improve Holodeck 2.0’s capacity for generating more complex and finely detailed 3D worlds, facilitating the creation of large-scale game environments. Additionally, we aim to integrate it with world models, enabling agents to navigate and interact with the generated environments.
[…]
Supplementary Material
6 Details of Holodeck 2.0
6.1 Efficiency Analysis
We evaluate the computational efficiency of Holodeck 2.0 on a single NVIDIA A6000 GPU. For a typical scene with 10–15 objects, the entire generation process takes about one hour, with 3D asset generation (via a locally deployed Hunyuan3D model) contributing roughly 80% of the runtime. The time scales approximately linearly with the number of objects, but parallelizing asset generation across multiple GPUs reduces total generation time to about 5 minutes.
6.2 Details of the Scene Analysis Module
@@ 12 I have the following list of image filenames:
6.3 Details of the Object Generation Module
In the Object Generation Module, we run the deployed Hunyuan3D 2.1 model locally. Given the individual image of each object, Hunyuan3D 2.1 generates a corresponding 3D asset in GLB-format. After generating the 3D assets for each object, we write the assets’ file path back into the JSON file containing the object properties. We also rescale each 3D asset based on its actual x, y, and z dimensions, using the original z size as a reference. The updated x and y sizes after scaling are then written back into the JSON file.
6.4 Details of the Scene Generation Module
@@
6.4.1 Spatial Constraints Definition
@@ 6.4.2 DFS Solver Implementation
3:  $\mathrm{best\_score}\leftarrow$-$\infty$, $\mathrm{best\_sol}\leftarrow\emptyset$
@@
6.5 Details of the Scene Editing Module
@@
6.6 Rendering Options
@@ 6.7 Gaming
7 Details of Evaluation
Notes on Baseline Coverage and Metric Availability.
When comparing Holodeck 2.0 with existing baselines, several practical constraints affect the completeness of certain metrics and the number of evaluated scenes.
First, GraphDreamer adopts an implicit 3D representation based on Gaussian splatting, where objects are not separable into distinct geometric units. Consequently, collision-based physical metrics (CF, CN, and PSA) cannot be computed, as they require explicit per-object geometry to detect intersections and evaluate physically grounded spatial consistency. In addition, the rendered views of GraphDreamer contain substantial visual clutter, making them unsuitable for human perceptual evaluation, which relies on clearly distinguishable objects and scene structure.
Second, GraphDreamer and Scene Language do not yield valid outputs for all scene descriptions in our benchmark. Among the 70 test scenes, GraphDreamer successfully produces only 14 scenes and Scene Language produces 44 scenes, whereas the other baselines and our method generate all 70. For fairness, we include all valid outputs returned by each method without imputing missing results.
These factors lead to different numbers of evaluated samples across baselines and explain why certain metrics are unavailable for methods whose representations or outputs do not permit their computation.
7.1 Human Evaluation
We conduct a total of four online questionnaires.
7.1.1 Comparative Analysis
The first two focus on scene generation quality in indoor and outdoor environments, respectively. Each questionnaire contains 35 distinct scenarios. For every scenario, participants answer eight rating questions on a 1–10 scale (1 = worst, 10 = best), evaluating:
  • Object-level quality across Holodeck, PartCrafter, Scene Language, and Holodeck 2.0
  • Scene-level quality across the above methods
These human-rated perceptual scores provide complementary evaluation to automated physical and spatial-semantic metrics, enabling a comprehensive comparison of all baselines against Holodeck 2.0.
@@ 7.1.1 Comparative Analysis
The rendered 3D scenes generated from this text are shown below.
(a) Scene A
(b) Scene B
(c) Scene C
(d) Scene D
@@
  5. 5.
    Please rate how well the objects generated in Scene C match the text input.
  6. 6.
    Please rate how well the overall scene in Scene C matches the text input.
  7. 7.
    Please rate how well the objects generated in Scene D match the text input.
  8. 8.
    Please rate how well the overall scene in Scene D matches the text input.
@@
The rendered 3D scenes generated from this text are shown below.
(a) Scene A
(b) Scene B
(c) Scene C
(d) Scene D
@@
  5. 5.
    Please rate how well the objects generated in Scene C match the text input.
  6. 6.
    Please rate how well the overall scene in Scene C matches the text input.
  7. 7.
    Please rate how well the objects generated in Scene D match the text input.
  8. 8.
    Please rate how well the overall scene in Scene D matches the text input.
7.1.2 Ablate the Scene Generation Module